Operations Report – 26 April 2018

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Status unchanged since yesterday.

Generator (hours run): 9:00 pm to 9:00 am.

Solar— SOC 80 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 305 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134126.7

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: Yes
Hours: 124.8
Beginning charge: 100%
Ending charge: 94 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.0
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: I took pans out of broiler underneath
oven; It is an older style gas stove, with a broiler (not a drawer)
underneath the oven. Too many pans there could block airflow for oven
burner.

Summary of internet: As of 9:30 AM, we have 40% of our internet
allowance remaining. The funny thing is that our internet while we
actually do have some percentage remaining doesn’t seem any faster
than at night when we are hard throttled.

Summary of suits and radios: All one piece suits are working (and
were used today). Radios are nominal.

Summary of Hab operations: Supported two (4 hour) EVAs.

Summary of GreenHab operations: Picked some plants for lunch today.

Summary of ScienceDome operations: Samples placed in incubator.

Summary of RAM operations: Morning EVA team brought materials in from
RAM upon returning. New radios are in the RAM airlock for us.

Summary of health and safety issues: None to report.